OpenStudy (anonymous):

H5P1 ZERO-OFFSET AMPLIFIER In many amplifiers, we use dual power supplies so we can obtain a 0V offset at the output, that is, the DC operating point at the output is 0V. An example amplifier circuit that can achieve this is shown below. In this circuit, VS+=1.0V, VS−=−1.0V, and the MOSFET parameters are K=1 mAV2 and VT=0.5V. what formulas do i need

OpenStudy (anonymous):

here i shall guide u through the solution , we have to find max Vin ; to find Vin let us apply kvll to the input loop Vin - Vgs+1 =0; ie, Vin = Vgs-1 ...............................(eqn1) now the unknown parameter is Vgs, hence to find it lets apply KVL to external loop : 1-Ids*R-Vds+1=0 , further substituting the value of Ids = k/2 *(Vgs-Vt)^2 in the above equation and solving the quadratic equation we will get, Vds = 0.593, -0.84 : as Vds needs tobe +ve hence neglecting he -ve term we have Vds = 0.593 and we know Vds = Vgs-Vt ie Vgs = Vds+Vt therefore, Vgs = 0.593+0.5=1.093 and substituting the above valuw in eqn1 we have Vin=1.093-1=0.093 #AMP#